OpenStudy (anonymous):

What is the distance between points a(6,-1) and b(-2,-1)?

OpenStudy (amistre64):

it the length of the hypotenuse created by ploting the points

OpenStudy (amistre64):



OpenStudy (amistre64):

or since its a straight line ....

OpenStudy (amistre64):

on an interval [a,b] the range, or distance, is equal to b-a

OpenStudy (anonymous):

\[d = \sqrt{(x_{a} - x_{b})^{2} + (y_{a}-y_{b})^{2}} = \sqrt{(6-(-2))^{2} + ((-1)-(-1))^{2}} = 8\]